OpenStudy (anonymous):

find and label the vortex and the line of symmetry. graf the function. f(x)=4x^2 the vertex is ?

OpenStudy (anonymous):

You find the vetex of a parabola using the formula -b/2a In your equation b=0 and a=4 so vertex is at (0,0) To find the axis of symmetry, you need to check the x-axis, y-axis, and origin. To check for symmetry around the x-axis, replace y with -y. If the equation remains unchanged then it is symmetrical around the x-axis. So pick a number for x, say 1 for example. Then y=4(1)^2 so y=4. Replacing y with -4 does not give the same value for x, so this equation in not symmetrical around the x- axis. To check for symmetry around th y-axis, replace x with a -x. If the equation remains unchanged then the equation is symmetrical around the y-axis. Using x=1 gives us y=4(1)^2 so y=4. Replacing x with -x gives us y=4(-1)^2 so y=4. So the equaton is symmetrical around the y-axis. To check for symmetry around the origin, replace x with -x and y with -y. If the equation remains the unchanged then it is symmetrical around the origin. So replacing x with -x and y with -y gives us -y=4(-x)^2. So for x=1 we have 4(-1)^2=-y so y=-4. So the equation is not symmetrical around the orgin.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

is there a way i can get you as a friend on face book or something so i can get more help latter this week?